OpenStudy (anonymous):

24 is 96% of what number?

OpenStudy (anonymous):

25...

OpenStudy (turingtest):

"24 is 96% of what number" can be written as\ 24=(0.96)x solve for x: x=24/0.96=? use a calculator!

OpenStudy (anonymous):

25

OpenStudy (anonymous):

25

OpenStudy (anonymous):

96x/100 = 24 x = 24 * 100/96 = 25